OpenStudy (christos):

Is my result of d/dx ln(ln(ln(x))) correct ? 1/((ln(ln(x)))^2*ln(ln(ln(x)))*ln(x)*x)

OpenStudy (anonymous):

very wierd problem ...

OpenStudy (anonymous):

well, d/dx ln( f(x) ) = 1/f(x) • f'(x) = f'(x)/f(x) So, that is the princple, so let's see...

OpenStudy (anonymous):

d/dx ln(ln(ln(x))) = 1/ln(ln(x)) • 1/(ln(x)) • 1/x So, d/dx ln(ln(ln(x))) = 1/\(\{\)x• ln(x) • ln[ln(x)]\(\}\) is my result

OpenStudy (inkyvoyd):

is it right? http://www.wolframalpha.com/input/?i=d%2Fdx+ln%28ln%28ln%28x%29%29%29 probably not

OpenStudy (anonymous):

yup, that is right inkyvoyd

OpenStudy (inkyvoyd):

I can do step by step if you want but you probably want to look over your work.. looks like you differentiated one iteration too many or the likes.

OpenStudy (anonymous):

and if you had d/dx \((\)ln(ln(ln x))\()^n\) then, you derivative would be d/dx (ln(ln(ln(x)))\(^n\) =\(\{\) n•ln(ln(ln(x)))\(^{n-1}\)\(\}\)\(/\)\(\{\){x• ln(x) • ln[ln(x)]\(\}\)

OpenStudy (anonymous):

I don't know how you got a ^2 in there tho

OpenStudy (anonymous):

you just do a chain rule like that, so then...

OpenStudy (anonymous):